OpenStudy (anonymous):

Solve: -2x - 13 = 8x + 7 -2 2 -6 6

OpenStudy (anonymous):

plz help

Nnesha (nnesha):

first subtract 8x both side let me know what you get

OpenStudy (anonymous):

wich side the 7 or 2

OpenStudy (anonymous):

do i subtract

Nnesha (nnesha):

yeah because you have to move 8x on other side or you cna add 2x both side instead subtracting 8x

OpenStudy (anonymous):

is it a negative number ?

Nnesha (nnesha):

which one

Nnesha (nnesha):

no 8x is positive

OpenStudy (anonymous):

-2 2 -6 6

OpenStudy (texaschic101):

what you do on one side of the equal sign, you have to do on the other

OpenStudy (anonymous):

i think the answer is a negativee

OpenStudy (anonymous):

but im not sure

OpenStudy (texaschic101):

subtract 8x from both sides...like this -2x - 13 = 8x + 7 --- subtract 8x from both sides -2x - 8x - 13 = 8x - 8x + 7 -- combine like terms -10x - 13 = 7 you understanding so far ? Then you would add 13 to both sides. why don't you try that and tell me what you get ?

OpenStudy (anonymous):

i got 2

Nnesha (nnesha):

positive 2??

OpenStudy (texaschic101):

-10x - 13 = 7 --- add 13 to both sides -10x - 13 + 13 = 7 + 13 --- combine like terms -10x = 20 -- divide both sides by -10 x = -2

Nnesha (nnesha):

when you move 13 to other side its become positive

OpenStudy (anonymous):

bc -10xdivided by10 =20 divided by 10 =2

OpenStudy (anonymous):

so is it b

OpenStudy (anonymous):

or a

OpenStudy (texaschic101):

its a....-2 because 20/-10 = -2

Nnesha (nnesha):

that suppose to -10

OpenStudy (texaschic101):

no...thats supposed to divide by -10

Nnesha (nnesha):

 after simplify this you get answer